# One-loop QCD corrections to SSA in unweighted Drell-Yan processes

### v1

@@ III Twist-3 distribution functions
$+H_{3}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5},\partial\psi}(k^{+},k_{1}^{+})+H_{4}^{\mu\nu}M^{(1)}_{\gamma^{+},\partial\psi}(k^{+},k_{1}^{+})+H_{5}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5},\partial G^{+}}(k^{+},k_{1}^{+})+H_{6}^{\mu\nu}M^{(1)}_{\gamma^{+},\partial G^{+}}(k^{+},k_{1}^{+})\Big{]}.$ (27)
@@
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}g_{0}^{\mu\nu}\times M^{(0)}_{\gamma^{+}\gamma_{5}\partial\psi}(k_{2}^{+})$
$+\int dk_{b}^{-}\bar{q}(x_{b})\int dk^{+}dk_{1}^{+}\Big{[}g_{1}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\partial\psi}+g_{2}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\gamma_{5}\partial\psi}+g_{3}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\partial_{\perp}G^{+}}+g_{4}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\gamma_{5}\partial_{\perp}G^{+}}\Big{]}$, (29)
@@
$s_{\perp}^{\rho}2\int dk^{+}P\frac{1}{k^{+}}\Big{(}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{)}$
$=$ $s_{\perp}^{\rho}M^{(0)}_{\gamma^{+}\gamma_{5},\partial\psi}(k_{2}^{+})-\int\frac{d\xi^{-}}{2\pi}e^{i\xi^{-}k_{2}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)(-i\gamma^{+}\gamma_{5})\partial_{\perp}^{\rho}\Big{(}\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})\Big{)}|ps\rangle+O((G^{+})^{2}).$ (36)
@@
$M^{(0)}_{\gamma^{+}\gamma_{5},\partial\psi}(k_{2}^{+})-2\int dk^{+}P\frac{1}{k^{+}}\Big{(}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big{)}=-2q_{\partial}(x_{2})+O((G^{+})^{2}).$ (37)
@@
After gauge link is recovered, remaining $G^{+}$ can be viewed as a part of gluon field strength tensor. We replace $\partial_{\perp}^{\rho}G^{+}$ in $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$ and $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$ into $G_{\perp}^{+\rho}$ as done in Qiu and Sterman (1999) or
@@
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}\tilde{g}_{0}^{\mu\nu}q_{\partial}(x_{2})$
$+\int dk_{b}^{-}\bar{q}(x_{b})\int dk^{+}dk_{1}^{+}\Big{[}\tilde{g}_{1}^{\mu\nu}M^{(1)}_{\gamma^{+}\partial\psi}+\tilde{g}_{2}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5}\partial\psi}-\frac{1}{\pi}\tilde{g}_{3}^{\mu\nu}T_{F}(x_{1},x_{2})-\frac{1}{\pi}\tilde{g}_{4}^{\mu\nu}T_{\Delta}(x_{1},x_{2})\Big{]},$ (40)
@@
$\tilde{g}_{0}=$ $-2g_{0},$
$\tilde{g}_{1}=$ $g_{1}-\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$
$\tilde{g}_{2}=$ $g_{2}+P\frac{2g_{0}(k_{2}^{+})}{k^{+}},$
$\tilde{g}_{3}=$ $g_{3}-\frac{1}{2}\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$
$\tilde{g}_{4}=$ $g_{4}+P\frac{2g_{0}(k_{2}^{+})}{k^{+}}.$ (41)
@@ IV Tree level result
with $\lambda$ a small quantity. $k$ and $k_{1}$ are also collinear to $p_{a}$, like $k_{2}$. At twist-3 level, the hard part should be expanded to $O(\lambda)$. One can find the details about collinear expansion in Qiu and Sterman (1999); Eguchi et al. (2007); Chen et al. (2017). Because the hard coefficients $H_{i}$ contain the delta function for momentum conservation $\delta^{n}(k_{2}+k_{b}-q)$, we also need to do power expansion for it, that is,
@@
$\partial\delta^{n-2}(q_{\perp})/\partial q_{\perp}^{\rho}$ gives the derivative part and $\delta^{(n-2)}(q_{\perp})$ gives the non-derivative part. Next we calculate these two parts separately.
@@
$H_{5}^{\mu\nu}|_{tree}=$ $H_{6}^{\mu\nu}|_{tree}=-4iN_{c}C_{F}\frac{1}{k_{b}\cdot k+i\epsilon}\frac{1}{p_{a}^{+}}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu}).$ (44)
@@
$H_{5}^{\mu\nu}|_{tree}=$ $H_{6}^{\mu\nu}|_{tree}=-4N_{c}C_{F}\delta(k_{b}\cdot k)\frac{1}{p_{a}^{+}}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu}).$ (46)
@@
$8N_{c}^{2}C_{F}W^{\mu\nu}_{non-de}=$ $-4\pi N_{c}C_{F}\int dk_{b}^{-}\int dk_{2}^{+}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})\Big{[}\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial G^{+}}(0,k_{2}^{+})\frac{(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu})}{p_{a}\cdot q}\Big{]}.$ (48)
@@
$H_{4}^{\mu\nu}=i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial k_{1\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}],\ H_{6}^{\mu\nu}=i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial k_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}].$ (49)
@@
$p_{a}^{+}H_{4}^{\mu\nu}|_{tree}=p_{a}^{+}H_{6}^{\mu\nu}|_{tree}=-i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}8N_{c}C_{F}\frac{1}{x+i\epsilon}g_{\perp}^{\mu\nu}.$ (50)
@@
In the last equality we have used eq.(39) to change $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$ to $T_{F}$. Above $W^{\mu\nu}_{tree}$ agrees with known result in Ma and Zhang (2015). It has been pointed out in Ma and Zhang (2015), such a structure of $W^{\mu\nu}$ with a derivative in $\delta^{(n-2)}(q_{\perp})$ satisfies QED gauge invariance in the sense of distribution, i.e.,
@@
with $t(q_{\perp})$ a test function which is normal at $q_{\perp}=0$.
@@ V Virtual corrections
In this section we present our results for one-loop virtual corrections. We first give the corrections to hadronic tensor $W^{\mu\nu}$. Same as tree level $W^{\mu\nu}$, the virtual correction contains nonderivative part and derivative part, which are calculated separately. Direct calculation of the one-loop integrals is complicated because a lot of tensor integrals are involved. A better method is to use FIRESmirnov (2008) to reduce these tensor integrals to standard scalar integrals. The reduced integrals are very simple: only standard two-point integrals remain. In the calculation, both UV and IR divergences are regulated by dimensional regularization, and we do not distinguish UV and IR divergences. UV divergences will be cancelled by counter term contributions discussed in Sec.7. After corrections to $W^{\mu\nu}$ are obtained, we give the result for $I\langle L\rangle$ and $I\langle P_{7}\rangle$.
@@ V.1 Corrections to nonderivative part of $W^{\mu\nu}$
Because $B(s_{2})-B_{c}(s_{2})$ is purely imaginary, only SGP gives nonzero contribution. Nonpole contribution is zero. $1/x_{1}$ is a PV, which is introduced by EOM relations. For $g_{2}$, we have
@@ V.1.1 SGP contribution
$s_{0}$ in right hand side eliminates the double pole $1/(x+i\epsilon)^{2}$ in $C_{3},D_{4}$. After this we get SGP contribution by the replacement $1/(x+i\epsilon)\rightarrow-i\pi\delta(x)$. The results are

### v2

@@ III Twist-3 distribution functions
$+H_{3}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+H_{4}^{\mu\nu}M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+H_{5}^{\mu\nu}M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})+H_{6}^{\mu\nu}M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big].$ (28)
@@
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}g_{0}^{\mu\nu}\times M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k_{2}^{+})$
$+\int dk_{b}^{-}\bar{q}(x_{b})\int dk^{+}dk_{1}^{+}\Big[g_{1}^{\mu\nu}\times M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}+g_{2}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}+g_{3}^{\mu\nu}\times M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}+g_{4}^{\mu\nu}\times M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}\Big$], (30)
@@
$2s_{\perp}^{\rho}\int dk^{+}P\frac{1}{k^{+}}\Big(M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big)$
$=$ $s_{\perp}^{\rho}M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k_{2}^{+})-\int\frac{d\xi^{-}}{2\pi}e^{i\xi^{-}k_{2}^{+}}\langle ps|\bar{\psi}(0)\mathcal{L}_{n}(0)(-i\gamma^{+}\gamma_{5})\partial_{\perp}^{\rho}\Big(\mathcal{L}_{n}^{\dagger}(\xi^{-})\psi(\xi^{-})\Big)|ps\rangle+O((G^{+})^{2}).$ (37)
@@
$M^{(0)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k_{2}^{+})-2\int dk^{+}P\frac{1}{k^{+}}\Big(M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}(k^{+},k_{1}^{+})\Big)=-2q_{\partial}(x_{2})+O((G^{+})^{2}).$ (38)
@@
After gauge link is recovered, remaining $G^{+}$ can be viewed as a part of gluon field strength tensor. We replace $\partial_{\perp}^{\rho}G^{+}$ in $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$ and $M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}G^{+}}$ into $G_{\perp}^{+\rho}$ as done in [4] or
@@
$8N_{c}^{2}C_{F}W^{\mu\nu}=$ $\int dk_{b}^{-}\bar{q}(x_{b})\int dk_{2}^{+}\tilde{g}_{0}^{\mu\nu}(k_{2}^{+})q_{\partial}(x_{2})$
$+\int dk_{b}^{-}\bar{q}(x_{b})\int dk^{+}dk_{1}^{+}\Big[\tilde{g}_{1}^{\mu\nu}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+},\partial_{\perp}\psi}(k^{+},k_{1}^{+})+\tilde{g}_{2}^{\mu\nu}(k^{+},k_{1}^{+})M^{(1)}_{\gamma^{+}\gamma_{5},\partial_{\perp}\psi}(k^{+},k_{1}^{+})$
$-\frac{1}{\pi}\tilde{g}_{3}^{\mu\nu}(k^{+},k_{1}^{+})T_{F}(x_{1},x_{2})-\frac{1}{\pi}\tilde{g}_{4}^{\mu\nu}(k^{+},k_{1}^{+})T_{\Delta}(x_{1},x_{2})\Big],$ (41)
@@
$\tilde{g}_{0}(k_{2}^{+})=$ $-2g_{0}(k_{2}^{+}),$
$\tilde{g}_{1}(k^{+},k_{1}^{+})=$ $g_{1}(k^{+},k_{1}^{+})-\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$
$\tilde{g}_{2}(k^{+},k_{1}^{+})=$ $g_{2}(k^{+},k_{1}^{+})+P\frac{2g_{0}(k_{2}^{+})}{k^{+}},$
$\tilde{g}_{3}(k^{+},k_{1}^{+})=$ $g_{3}(k^{+},k_{1}^{+})-\frac{1}{2}\delta(k^{+})g_{1}^{SGP}(k_{1}^{+}),$
$\tilde{g}_{4}(k^{+},k_{1}^{+})=$ $g_{4}(k^{+},k_{1}^{+})+P\frac{2g_{0}(k_{2}^{+})}{k^{+}}.$ (42)
@@ IV Tree level result
with $\lambda$ a small quantity. $k$ and $k_{1}$ are also collinear to $p_{a}$, like $k_{2}$. At twist-3 level, the hard part should be expanded to $O(\lambda)$. One can find the details about collinear expansion in [4, 7, 15]. Because the hard coefficients $H_{i}$ contain the delta function for momentum conservation $\delta^{n}(k_{2}+k_{b}-q)$, we also need to do power expansion for it, that is,
@@
$\partial\delta^{n-2}(q_{\perp})/\partial q_{\perp}^{\rho}$ gives the derivative part and $\delta^{(n-2)}(q_{\perp})$ gives the non-derivative part. Such an expansion is necessary because we have to integrate out $q_{\perp}$ in the cross section. For example, considering a test function $t(q_{\perp})$, we have
$\int d^{n-2}q_{\perp}t(q_{\perp})\delta^{(n-2)}(q_{\perp}-k_{2\perp})=t(k_{2\perp}).$ (45)
Then, collinear expansion gives
$t(k_{2\perp})=t(0)+\frac{\partial t(k_{2\perp})}{\partial k_{2\perp}^{\rho}}\Big|_{k_{2\perp}=0}k_{2\perp}^{\rho}+O(k_{2\perp}^{2}).$ (46)
Now the first and the second terms on RHS can be reproduced by the non-derivative and derivative terms of the expansion of $\delta^{(n-2)}(q_{\perp}-k_{2\perp})$. Next we calculate these two parts separately.
@@
$H_{5}^{\mu\nu}|_{\text{tree}}^{\text{Fig.1b}}=$ $H_{6}^{\mu\nu}|_{\text{tree}}^{\text{Fig.1b}}=-2iN_{c}C_{F}\frac{1}{k_{b}\cdot k+i\epsilon}\frac{1}{p_{a}^{+}}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu}).$ (47)
@@
$H_{5}^{\mu\nu}|_{\text{tree}}^{\text{Fig.1b+c.c}}=$ $H_{6}^{\mu\nu}|_{\text{tree}}^{\text{Fig.1b+c.c}}=-4\pi N_{c}C_{F}\delta(k_{b}\cdot k)\frac{1}{p_{a}^{+}}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu}).$ (49)
@@
$8N_{c}^{2}C_{F}W^{\mu\nu}_{non-de}=$ $-4\pi N_{c}C_{F}\int dk_{b}^{-}\int dk_{2}^{+}\delta(k_{2}^{+}-q^{+})\delta(k_{b}^{-}-q^{-})\delta^{n-2}(q_{\perp})\Big[\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}(0,k_{2}^{+})\frac{(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu})}{p_{a}\cdot q}\Big].$ (51)
@@
$H_{4}^{\mu\nu}=$ $i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial k_{1\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}\otimes\gamma^{+}],$
$H_{6}^{\mu\nu}=$ $i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial k_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}Tr[H^{\mu\nu}\otimes T^{a}\gamma^{-}\otimes\gamma^{+}].$ (52)
@@
$p_{a}^{+}H_{4}^{\mu\nu}|_{\text{tree}}^{\text{Fig.1b}}=p_{a}^{+}H_{6}^{\mu\nu}|_{\text{tree}}^{\text{Fig.1b}}=-i\frac{\partial\delta^{n}(k+k_{1}+k_{b}-q)}{\partial q_{\perp}^{\rho}}\tilde{s}_{\perp}^{\rho}4N_{c}C_{F}\frac{1}{x+i\epsilon}g_{\perp}^{\mu\nu}.$ (53)
@@
In the last equality we have used eq.(40) to change $M^{(1)}_{\gamma^{+},\partial_{\perp}G^{+}}$ to $T_{F}$. Above $W^{\mu\nu}_{tree}$ agrees with known result in [17]. It has been pointed out in [17], such a structure of $W^{\mu\nu}$ with a derivative in $\delta^{(n-2)}(q_{\perp})$ satisfies QED gauge invariance in the sense of distribution, i.e.,
@@
with $t(q_{\perp})$ a test function which is normal at $q_{\perp}=0$. With explicit expression of lepton momentum $l^{\mu}$ given in eq.(8), we get
$(p_{a}^{\mu}\tilde{s}_{\perp}^{\nu}+p_{a}^{\nu}\tilde{s}_{\perp}^{\mu})L_{\mu\nu}\Big|_{q_{\perp}=0}=\frac{4s_{2}}{x_{2}}\cos\theta\tilde{s}_{\perp}\cdot l_{\perp,cs},\quad g_{\perp}^{\mu\nu}\frac{\partial L_{\mu\nu}}{\partial q_{\perp}^{\rho}}\Big|_{q_{\perp}=0}=0.$ (58)
The second equation indicates that derivative term does not contribute to SSA. We thus get the asymmetry
$A_{UT}\equiv$ $\frac{\frac{d\sigma(s_{\perp})}{dQ^{2}d\Omega}-\frac{d\sigma(-s_{\perp})}{dQ^{2}d\Omega}}{\frac{d\sigma(s_{\perp})}{dQ^{2}d\Omega}+\frac{d\sigma(-s_{\perp})}{dQ^{2}d\Omega}}=-\frac{\sin 2\theta\sin\phi}{2Q(1+\cos^{2}\theta)}\frac{\int\frac{d\xi}{\xi}\frac{dx_{b}}{x_{b}}\frac{dx_{2}}{x_{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})\bar{q}(x_{b})T_{F}(x_{2},x_{2})}{\int\frac{d\xi}{\xi}\frac{dx_{b}}{x_{b}}\frac{dx_{2}}{x_{2}}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})\bar{q}(x_{b})q(x_{2})}.$ (59)
Chiral-Odd contribution is not included. As illustrated in [17], this is the same as [29], but only $1/2$ of [30, 31, 32], [33]. We also notice that if one ignores the time-reversal odd function in [30, 31, 32], i.e., $f_{T}(x)=0$ in their notation, the result becomes the same. $f_{T}(x)=0$ is equivalent to our $M^{(0)}_{\gamma_{\perp}}=0$ as indicated by our eq.(21). This equation is a consequence of PT symmetry. Since we work in Feynman gauge, for which gauge potentials vanish at infinity, we think PT symmetry is preserved and this equation is reasonable. We have used this equation in our calculation. One can find further discussions about this discrepancy in [17].
@@ V Virtual corrections
In this section we present our results for one-loop virtual corrections. We first give the corrections to hadronic tensor $W^{\mu\nu}$. Same as tree level $W^{\mu\nu}$, the virtual correction contains nonderivative part and derivative part, which are calculated separately. Direct calculation of the one-loop integrals is complicated because a lot of tensor integrals are involved. A better method is to use FIRE[34] to reduce these tensor integrals to standard scalar integrals. The reduced integrals are very simple: only standard two-point integrals remain. In the calculation, both UV and IR divergences are regulated by dimensional regularization, and we do not distinguish UV and IR divergences. UV divergences will be cancelled by counter term contributions discussed in Sec.VII. After corrections to $W^{\mu\nu}$ are obtained, we give the result for $I\langle L\rangle$ and $I\langle P_{7}\rangle$.
@@ V.1 Corrections to nonderivative part of $W^{\mu\nu}$
Because $B(s_{2})-B_{c}(s_{2})$ is purely imaginary, only SGP gives nonzero contribution. $1/x_{1}$ is a PV, which is introduced by EOM relations. For $g_{2}$, we have
@@ V.1.1 SGP contribution
$s_{0}$ on right hand side eliminates the double pole $1/(x+i\epsilon)^{2}$ in $C_{3},D_{4}$. After this we get SGP contribution by the replacement $1/(x+i\epsilon)\rightarrow-i\pi\delta(x)$. The results are